OpenStudy (anonymous):

Please help!! This is on Two-step equations ill post the question in the comments because it wont fit

OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $14,000. Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $2,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 2% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time. Add the amount of interest to the price of the car.) Which monthly payment amount is lower? Please explain how you arrived at your answer and show all of your work.

OpenStudy (anonymous):

Option 1: Amount financed = purchase price - down payment = 14000 - 2000 = 12000 Add on interest = principal * rate * time = 0 Monthly payments: (amount financed + add on interest)/number of months = 12000/60 = 200 Option 2: Amount financed = 14000 Add on interest = 14000*.02*6 = 1680 Monthly payments: (14000+1680)/72 = 217.78

OpenStudy (anonymous):

so is it the first one? >.<

OpenStudy (anonymous):

I have option 1 and option 2

OpenStudy (anonymous):

yea but it lastly asks Which monthly payment amount is lower? Please explain how you arrived at your answer and show all of your work.

OpenStudy (anonymous):

if you look you will see the answer

OpenStudy (anonymous):

so the first one because the amout is lower right ?

OpenStudy (anonymous):

*amount

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem